Korean VLBI Network (KVN)

Observation status

Return to List

Exp_code:  p15sl01f Season  2015B
Title:  iMOGABA25 (reassigned) PI: Name: Sang-Sung Lee, E-mail: sslee@kasi.re.kr
UT Time: 2015-08-28 02:00 ~ 2015-08-29 02:00 Band:  K22GHz,Q43GHz,W86GHz,D129GHz
Polarization: LHCP Data rate: 1Gbps
Observation comment
Observation status End of observation.
Observation summary view
CodeProblemsKYSKUSKTN
120Problems with recorder1
130Severe weather condition1
200DAS & Clock errors1
  Summary resultSMS
S:Success, M:Success with minor failures, F:Failure, C:Cancel
Correlation
Media arrival date:  Media Copy date  2015-09-18
Correlation date:  2015-09-22 Release date 
Public open date:  Fits file: 
Correlation Comment:
YS, TN : 1Gbps
US : 2Gbps

Fringe

Filename Result Comment
 

Station report

KYS station
Exp_code  p15sl01f Obs Time(UT):  2015-08-28 02:00 ~ 2015-08-29 02:00
result:  Success Weather:  Cloudy(26.6C, 62.5%, 0.0m/s, 998.6mbar)
Clock Offset:  -0.3 Diskpack:  KVN-2105/16000/1024
Pointing:  (-3, 1) Focus:  400
Comment:  Scan #19(UT 04:33) - Scan #21(UT 04:55) :Sometimes Rain
Log file:   p15sl01fKYS.log
Antab file:   p15sl01fKYS.ANTAB.1
Antab image:  p15sl01fKYS.png
KUS station
Exp_code  p15sl01f Obs Time(UT):  2015-08-28 02:00 ~ 2015-08-29 02:00
result:  Success with minor failures Weather:  Clear(26.2C, 59.9%, 2.5m/s, 995.0mbar)
Clock Offset:  -0.5 Diskpack:  KVN-3179/24000/1024
Pointing:  (-2.1, 8.7) Focus:  -210
Comment:  scan #1; Recorder overflow error-> recording to No0001a(UT 02:05)
Scan #42~46 (UT 07:50 ~ 08:25) : Rain
Scan #87~#93 (UT 14:03-14:56): no response from Rx-> Rx restart
no response occur again -> Das, Rx restart
Problems:  1: 2015-08-28 02:00~02:05 Problems with recorder: Recorder overflow error-> recording to No0001a
2: 2015-08-28 07:50~08:25 Severe weather condition: Rain
3: 2015-08-28 14:03~14:56 DAS & Clock errors: no response from Rx-> Rx restart but occur again -> Das, Rx restart
Log file:   p15sl01fKUS.log
Antab file:   p15sl01fKUS.ANTAB.1
Antab image:  p15sl01fKUS.png
KTN station
Exp_code  p15sl01f Obs Time(UT):  2015-08-28 02:00 ~ 2015-08-29 02:00
result:  Success Weather:  Clear(26.5C, 77.3%, 1.1m/s, 963.5mbar)
Clock Offset:  0.2 Diskpack:  KVN-3150/24000/1024
Pointing:  (-5.1, 1.7) Focus:  140
Comment: 
Log file:   p15sl01fKTN.log
Antab file:   p15sl01fKTN.ANTAB.1
Antab image:  p15sl01fKTN.png

KASI logo image

Korean VLBI Network